OpenStudy (anonymous):

The size of a certain cell is 2.5 × 10^–9 m. Another cell is 1.5 × 10^3 times larger. How large is the larger cell in scientific notation? a.)4 × 10^–27m b.)4 × 10^–12m c.)3.75 × 10^6m d.)3.75 × 10^–6m

OpenStudy (anonymous):

hello sister.. answer ish--> 5. (2.5 × 10^–9)(1.5 × 10^3) = 2.5*1.5 x 10^(-9 + 3) = 3.75 x 10^-6 m

OpenStudy (anonymous):

Thank you sir'(:

OpenStudy (anonymous):

am a 5 year old kid with 10 years of experience :-)